A parabola "y=ax^{2}+bx+c"  passes through the points (-2, 10); (1,4) and (2,6).

Which of the following are correct? More than one answer may be correct


  1. The augmented matrix for the system of equations connecting a, b and c is "\left(\begin{array}{cccc}4&-2&1&:&10\\1&1&1&:&4\\4&2&1&: &6\end{array}\right)"
  2. The determinant of the coefficient matrix for the system of equations connecting a, b and c is -12.
  3. The parabola above passes through the origin.
  4. By applying Gauss elimination, the normal form of the augmented matrix for the system of equations connecting a, b and c is"\\left(\\begin{array}{cccc}1&0&0&:&1\\\\0&1&0&:&-1\\\\0&0&1&: &4\\end{array}\\right).\n\n\u200b"
1
Expert's answer
2021-08-31T13:20:32-0400

"(-2, 10)"


"a(-2)^2+b(-2)+c=10"


"(1,4)"


"a(1)^2+b(1)+c=4"

"(2,6)"


"a(2)^2+b(2)+c=6"


"\begin{matrix}
4a-2b+c=10\\
a+b+c=4\\
4a+2b+c=6\\
\end{matrix}"

1. The augmented matrix for the system of equations connecting a, b and c is


"A=\begin{pmatrix}
 4 & -2 & 1 & : & 10 \\
 1 & 1 & 1 & : & 4 \\
 4 & 2 & 1 & : & 6 \\
\end{pmatrix}"

True.


2.


"\begin{vmatrix}
 4 & -2 & 1 \\
 1 & 1 & 1 \\
 4 & 2 & 1 \\
\end{vmatrix}=4\begin{vmatrix}
 1 & 1 \\
 2 & 1
\end{vmatrix}-(-2)\begin{vmatrix}
 1 & 1 \\
 4 & 1
\end{vmatrix}+1\begin{vmatrix}
 1 & 1 \\
 4 & 2
\end{vmatrix}"


"=-4-6-2=-12"

True.


3.

"(0,0)"


"a(0)^2+b(0)+c=0=>c=0"


"\begin{matrix}
4a-2b+0=10\\
a+b+0=4\\
4a+2b+0=6\\
\end{matrix}"




"\begin{matrix}
8a=16\\
a+b=4\\
4a+2b=6\\
\end{matrix}"




"\begin{matrix}
a=2\\
b=2\\
4a+2b=6\\
\end{matrix}"


No solution.


False.


4.


"A=\begin{pmatrix}
 4 & -2 & 1 & : & 10 \\
 1 & 1 & 1 & : & 4 \\
 4 & 2 & 1 & : & 6 \\
\end{pmatrix}"

"R_1=R_1\/4"


"\begin{pmatrix}
 1 & -1/2 & 1/4 & : & 5/2 \\
 1 & 1 & 1 & : & 4 \\
 4 & 2 & 1 & : & 6 \\
\end{pmatrix}"

"R_2=R_2-R_1"


"\begin{pmatrix}
 1 & -1/2 & 1/4 & : & 5/2 \\
 0 & 3/2 & 3/4 & : & 3/2 \\
 4 & 2 & 1 & : & 6 \\
\end{pmatrix}"

"R_3=R_3-4R_1"


"\begin{pmatrix}
 1 & -1/2 & 1/4 & : & 5/2 \\
 0 & 3/2 & 3/4 & : & 3/2 \\
 0 & 4 & 0 & : & -4 \\
\end{pmatrix}"

"R_2=2R_2\/3"


"\begin{pmatrix}
 1 & -1/2 & 1/4 & : & 5/2 \\
 0 & 1 & 1/2 & : & 1 \\
 0 & 4 & 0 & : & -4 \\
\end{pmatrix}"

"R_1=R_1+R_2\/2"


"\begin{pmatrix}
 1 & 0 & 1/2 & : & 3 \\
 0 & 1 & 1/2 & : & 1 \\
 0 & 4 & 0 & : & -4 \\
\end{pmatrix}"

"R_3=R_3-4R_2"


"\begin{pmatrix}
 1 & 0 & 1/2 & : & 3 \\
 0 & 1 & 1/2 & : & 1 \\
 0 & 0 & -2 & : & -8 \\
\end{pmatrix}"

"R_3=-R_3\/2"


"\begin{pmatrix}
 1 & 0 & 1/2 & : & 3 \\
 0 & 1 & 1/2 & : & 1 \\
 0 & 0 & 1 & : & 4 \\
\end{pmatrix}"

"R_1=R_1-R_3\/2"


"\begin{pmatrix}
 1 & 0 & 0 & : & 1 \\
 0 & 1 & 1/2 & : & 1 \\
 0 & 0 & 1 & : & 4 \\
\end{pmatrix}"

"R_2=R_2-R_3\/2"


"\begin{pmatrix}
 1 & 0 & 0 & : & 1 \\
 0 & 1 & 0 & : & -1 \\
 0 & 0 & 1 & : & 4 \\
\end{pmatrix}"

The normal form of the augmented matrix for the system of equations connecting a, b and c is


"\begin{pmatrix}
 1 & 0 & 0 & : & 1 \\
 0 & 1 & 0 & : & -1 \\
 0 & 0 & 1 & : & 4 \\
\end{pmatrix}"

True.